OpenStudy (anonymous):

how is agrpagh suppose to look like with this equation? correctly represents the equation 3x + 4y = 12.

OpenStudy (anonymous):

What is the equation of the line that passes through the points (−2, 1) and (1, 10)? 3x − y = −7 3x − y = −5 3x − y = 5 x + 3y = −5